Sin(75°) Kskkjsjsjssksoeihevv hdjdjdjjfjdhdvd f fbfjfjdjkoekehevvebdhdjrb rjfjfbr fbfjfifhdbjddivej99393838389398384838377n dd dj8383838838484uudjbbsekj(+|++|*|*|+#&#*#×)jishsjishdhduddieeihsj8gwjibb2 bejejjbssoeeujeiieeiieehbeheie9e8 eueiehdeoeigebwjwiegejejeiue ebieirhrb3i9e8393937383983638383737338374848848484yevebheie8dhdhd. Dhdidd r dJrirbr rbrjdr. Dibrbrjd r ridnr r didjdnd d. D.jdidnd r fjfjrbr r rbrjf

Sin(75°)
Kskkjsjsjssksoeihevv hdjdjdjjfjdhdvd f fbfjfjdjkoekehevvebdhdjrb rjfjfbr fbfjfifhdbjddivej99393838389398384838377n dd dj8383838838484uudjbbsekj(+|++|*|*|+#&#*#×)jishsjishdhduddieeihsj8gwjibb2 bejejjbssoeeujeiieeiieehbeheie9e8 eueiehdeoeigebwjwiegejejeiue ebieirhrb3i9e8393937383983638383737338374848848484yevebheie8dhdhd. Dhdidd r dJrirbr rbrjdr. Dibrbrjd r ridnr r didjdnd d. D.jdidnd r fjfjrbr r rbrjf 

Grade:11

1 Answers

Aditya Gupta
4 years ago
hello rachit. in think you mean to find the value of sin75.
sin75= sin(45+30)= sin45*cos30 + sin30*cos45
= (1/√2)*√3/2 + (1/2)*(1/√2)
(1/2√2)*(√3 + 1)
(√3 + 1)/2√2
= (√6 + √2)/4
kindly approve :)